Kok Kiang Company buys a bus at RM140,000. The useful life of the bus is eight years. If the book value at the end of four years is RM80,000, find the scrap value of the bus at the end of eight years using the straight line method.


1
Expert's answer
2021-01-19T17:48:02-0500

Answer:

Given-

Cost of bus= RM140,000

Useful life of the bus= 8 years

Book value at the end of four years= RM80,000

Scrap value of the bus at the end of eight years using SLM=?


Depreciated amount= 140,000 - 80,000 = RM60,000

RM60,000 more will be depreciated in next 4 years

So, the scrap value of the bus at the end of eight years is 80,000 - 60,000 = RM20,000